OpenStudy (anonymous):

Okay, so I have a few problems to do, but I have no clue how to solve them. It is geometry proofs using the distance formula I will post the pictures here

OpenStudy (anonymous):

there are the first 3 problems. If anyone is patient enough to stay and help me with them until I get it would be greatly appreciated

OpenStudy (bibby):

do you know the formula to find slope given 2 points?

OpenStudy (anonymous):

y2-y1 ------= slope? x2-x1

OpenStudy (bibby):

looks right. how would you go about finding the slope of AC?

OpenStudy (anonymous):

A(0,0) C(a,a) a-0 ----= ? a-0

OpenStudy (bibby):

1

OpenStudy (bibby):

you can simplify it because a-0=a and a/a=1

OpenStudy (anonymous):

okay, I get it that part.

OpenStudy (bibby):

try bd

OpenStudy (anonymous):

B(a,0) D(0,a) a-0 a ---- = --= -1? 0-a -a

OpenStudy (bibby):

correct. what do you know about the slopes of perpendicular lines?

OpenStudy (anonymous):

You have to take the reciprocal of the original slope. Like 1/3 would be -3/1 or just -3

OpenStudy (bibby):

exactly. so we have slopes 1 and -1 of AC/BD respectively

OpenStudy (bibby):

 what do es that mean

OpenStudy (anonymous):

Line AC is perpendicular to line ___ would it be BD

OpenStudy (bibby):

yep

OpenStudy (anonymous):

wow, okay that was easy. Now it's the other ones that scare me haha

OpenStudy (bibby):

the 2nd one looks bleh :/ let's start with P use the fact that the midpoint for \((x_1,y_1) \text {and}(x_2,y_2)\) is \((\dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2})\)

OpenStudy (anonymous):

okay, I know how to solve for midpoint but p has a two in it and I'm not sure where that is coming from

OpenStudy (bibby):

what do you mean?

OpenStudy (anonymous):

P= 0+2a\2 Where is the 2a part from on the diagram?

OpenStudy (bibby):

P is the midpoint of A and B, right? so our x coordinate is half (you can think of it as the average) of the sum of the 2 x coordinates 0+2a

OpenStudy (bibby):

the 0 is the x coord of A the 2a is the x coord of B

OpenStudy (anonymous):

ooohhh, okay! so they're meeting in the middle of P I was thinking it was using the points of P. Makes so much more sense now.

OpenStudy (bibby):

do you want to fill in the rest of the stuff for P?

OpenStudy (bibby):

or does anything else need explaining

OpenStudy (anonymous):

I need to kind of be walked through it. This thing online is so confusing so when someone is explaining it me it makes more sense.

OpenStudy (anonymous):

i'll try writing it out then show it to you

OpenStudy (bibby):

cool, cool that works

OpenStudy (anonymous):